• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 17:07:41 on June30,2010

    Entry number 329213

    This entry is a followup to: 329198

    Followups:

    keyword=trig / daq status update

    With a lot of work done by Neil and Kurtis, here's the status

    The pulser goes on a cable from R-HRS to L-HRS which is 232 nsec
    long, measured with reflections. The pulser on R-HRS is compensated
    so the signals arrive at the S2m electronics at the same time.

    Coinc of S2m on L and R-HRS:
    The overlap at the coincidence circuit is good. Two signals 10 nsec
    wide with the R-HRS late by 3 nsec.

    The delay of T1 to TS (trig. super.) did not need to be adjusted,
    though there was a slight complication that we had to bypass the
    old route. So the gates on L-HRS should be good and unchanged.
    The R-HRS trig. and coinc. trigger were lined up to match T1 at TS.

    The triggers are

    T1 = L-HRS S2m
    T2 = L-HRS S0
    T3 = R-HRS S2m
    T4 = coincidence of T1 and T3
    T5 = L-HRS singles from a coinc of S0 and GasCer.
    T6 = T4.and.GasCer on R-HRS (added later, see halog 329311).
    T8 = 1024 Hz pulser

    T1, T2, T5, and T8 were not changed from previous running.

    A "to do" list which I think takes 5 hours:

    1. Some fine-tuning of the alignment of timing at TS.
    2. Scope checks of gate timing (repeat L-HRS, do the R-HRS 1st time)
    3. Put all triggers into a TDC to measure coinc. time.
    4. Use front-panel gates for ADCs on R-HRS (presently back-panel).
    5. Finish scaler cabling.
    6. Formation of the coinc trigger S2m-L.and.S2m-R.and.GasCer-R
    7. Fix / replace the TDC in ROC1 slot 11 (separate halog entry).

    I think the timing on R-HRS may be good enough for VDC checkout.